1) A sample of PVC pipes coming off a production line were tested for pipe diameter size. The statistical results (in millimeters) were:
Mean 300
Median 300
Mode 300
Standard Deviation 15
Range 90
Number in Sample 100
a) According to the Normal Rule, what percent of the pipes had diameters between 285 and 315?
b) What would the range of pipe diameters need to be in order to capture 95% of the pipe diameters?

Answer by stanbon(48568)

1) A sample of PVC pipes coming off a production line were tested for pipe diameter size. The statistical results (in millimeters) were:
Mean 300
Median 300
Mode 300
Standard Deviation 15
Range 90
Number in Sample 100
a) According to the Normal Rule, what percent of the pipes had diameters between 285 and 315?
Find z-scores for 285 and for 315:
z(285) = (285-300)/15=-1
z(315) = (315-300)/15=+1
------------
Then P(-1 ============================
b) What would the range of pipe diameters need to be in order to capture 95% of the pipe diameters?
Find the z-score corresponding to 95%
Use your z-chart or InvNorm(0.95) to get z=1.645
---------
Then find the raw score associated with that z-score:
1.645 = (x-300)/15
x-300 = 24.67
x = 324.67
=====================
Cheers,
Stan H.